# Protecting multimodal large language models against misleading visualizations

### v4

@@ Abstract
Visualizations play a pivotal role in daily communication in an increasingly data-driven world. Research on multimodal large language models (MLLMs) for automated chart understanding has accelerated massively, with steady improvements on standard benchmarks. However, for MLLMs to be reliable, they must be robust to misleading visualizations, i.e., charts that distort the underlying data, leading readers to draw inaccurate conclusions that may support disinformation. Here, we uncover an important vulnerability: MLLM question-answering (QA) accuracy on misleading visualizations drops on average to the level of the random baseline. To address this, we introduce the first inference-time methods to improve QA performance on misleading visualizations, without compromising accuracy on non-misleading ones. We find that two methods, table-based QA and redrawing the visualization, are effective, with improvements of up to 19.6 percentage points. We make our code and data available.111github.com/UKPLab/arxiv2025-misleading-visualizations
@@ 1 Introduction
In an increasingly data-driven world, visualizations are widely used by scientists, journalists, governments, or companies to efficiently communicate data insights to a broad audience (Huang et al., 2025). They play a crucial role in crisis settings, such as during the COVID-19 pandemic, where charts were shared daily to inform the public about the spread of the virus (Zhang et al., 2021). In many cases, visualizations support a message more convincingly than if the underlying data table was shown directly to readers (Pandey et al., 2014). However, visualizations can also be misleading (Correll and Heer, 2017). This occurs when design flaws, or misleaders, distort the underlying data in a way that prevents a correct interpretation by the reader (Tufte and Graves-Morris, 1983; Pandey et al., 2015; Correll and Heer, 2017; Lauer and O’Brien, 2020; McNutt et al., 2020; Yang et al., 2021; Lo et al., 2022; Lisnic et al., 2023; Lan and Liu, 2025). These misleaders appear across a wide range of visualizations, including bar, pie, and line charts, choropleth maps, or scatter plots. Recent taxonomies have documented over 70 distinct types of misleaders observed in real-world examples (Lo et al., 2022; Lan and Liu, 2025). Major categories include axes manipulations, such as truncated and inverted axes, or the use of visual manipulations like area encoding and 3D effects. Figure 1 shows three real-world examples of misleading visualizations (Lo et al., 2022) paired with multiple-choice questions. Given a question about the underlying data table, their misleaders lead readers to infer a misleading answer instead of the correct one. Figure 2 illustrates how two visualizations of the same underlying data, one non-misleading and the other misleading, can lead to different interpretations by readers.
Figure 1: Three examples of real-world misleading visualizations (Lo et al., 2022) with multiple choice questions. The correct answer is colored in green, while the wrong answer supported by the misleader is colored in purple.
Figure 2: Non-misleading and misleading visualizations of the same data (Lauer and O’Brien, 2020) with a Likert-scale question where 1 means “a little” and 6 means “a lot”. A consistent interpretation requires identical responses for both visualizations. However, the truncated axis deceives readers in choosing a higher value for the misleading visualization.
Prior work has revealed the potential of misleading visualizations to deceive human readers. Some studies (Pandey et al., 2015; Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024) have shown that such visualizations reduce human accuracy when answering multiple-choice, free-text, or rank questions about the underlying data. Other studies (Pandey et al., 2015; O’Brien and Lauer, 2018; Lauer and O’Brien, 2020) have demonstrated that human readers tend to provide different Likert-scale answers to a question depending on whether they view a misleading or non-misleading visualization of the same data table.
@@
Multimodal large language models (MLLMs) have sparked a massive interest in automated chart understanding research (Huang et al., 2025), with steady progress on reference benchmarks such as ChartQA (Masry et al., 2022; Hoque et al., 2022), highlighting their strong potential as interactive assistants for extracting data insights. However, existing research has largely overlooked the important threat posed by misleading visualizations. If MLLMs, like humans, are easily deceived by misleading visualizations, malicious actors could exploit this vulnerability. MLLMs confronted to misleading visualizations could present incorrect interpretations of the underlying data to users, contributing to the spread of disinformation, and reinforcing human belief in it.
In this work, we conduct the first comprehensive study to assess and mitigate the vulnerability of 18 MLLMs of varying sizes to misleading visualizations. In Experiment 1, we compare the question-answering (QA) accuracy of MLLMs on misleading and non-misleading visualizations. In Experiment 2, we evaluate the consistency of MLLMs, i.e., whether they provide the same response to a Likert-scale question when shown either a misleading or a non-misleading visualization of the same underlying data. The results of both experiments demonstrate that MLLMs, like humans, are indeed vulnerable to misleading visualizations. In Experiment 3, we introduce six inference-time correction methods to reduce the vulnerability of MLLMs to misleading visualizations in a QA setting. We evaluate them on the datasets of Experiment 1. The goal is to improve QA accuracy on misleading visualizations while maintaining performance on non-misleading ones.
In summary, our contributions are twofold: (1) We present the first extensive analysis of the vulnerabilities of 18 MLLMs to 17 types of misleading visualizations, evaluating both QA accuracy and Likert-scale consistency. (2) We are the first to introduce six correction methods to mitigate the negative impact of misleaders on QA tasks.
@@ 2.1 Datasets
Experiments 1 and 3 We compare the question-answering (QA) accuracy of MLLMs across three datasets: (a) a misleading visualization dataset containing $n$ = 143 instances, and featuring 17 types of misleaders, defined in Table 2 (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024; Lo et al., 2022); (b) a non-misleading visualization dataset ($n$ = 124) (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024; Lee et al., 2017), also combining real-world and synthetic cases; and (c) ChartQA (Masry et al., 2022), the reference real-world non-misleading benchmark ($n$ = 2500). Datasets (a) and (b) combine three existing resources and one introduced in this work. First, CALVI (Ge et al., 2023) includes 45 misleading and 15 non-misleading visualizations based on synthetic data, each paired with a multiple-choice question (MCQ). The dataset was initially designed to evaluate humans. Second, CHARTOM (Rho et al., 2023; Bharti et al., 2024) contains 56 samples, including 28 MCQs, 20 free-text questions, and 8 rank questions. Each question is linked to two visualizations, one misleading and one non-misleading. Like CALVI, the underlying data is synthetic, and the test was originally designed to evaluate humans. Third, VLAT (Lee et al., 2017), the reference dataset to evaluate humans, provides 12 non-misleading visualizations, each paired with three to seven MCQs, for a total of 53 instances. The visualizations are based on real-world data. Fourth, we introduce a dataset of 42 real-world misleading visualizations, each annotated with a MCQ with three to four choices. They come from a collection annotated with the misleaders affecting them (Lo et al., 2022), which inspired the synthetic examples in CALVI. We manually create MCQs, using as reference those from CALVI and CHARTOM. By incorporating visualizations and questions about real-world data, we introduce direct conflicts with MLLMs’ parametric knowledge, allowing us to assess their vulnerability in real-world scenarios.
Experiment 2 We use four pairs of visualizations, one misleading and the other non-misleading, originally designed by Lauer and O’Brien (2020) to evaluate humans. The pairs cover three distinct misleaders.
@@ 2.2 Correction methods
Providing (2) the axes data, (3) the data table, (4) or both: we prompt the MLLM in zero-shot to extract the axes labels or underlying data table. The axes and tables are formatted as text strings. We do not impose constraints on the delimiters used to indicate new rows and columns. The axes, the table, or both are then provided as additional input in the prompt.
(5) Table-based QA: after extracting the table with the MLLM, we provide it alone to a text-only LLM, reframing the task as table-based QA.
(6) Redrawn visualization: after extracting the table with the MLLM, we provide it to a text-only LLM which generates Python code to create a visualization using Matplotlib (Hunter, 2007). If the code compiles successfully, the generated visualization replaces the original one in the QA prompt; otherwise, the original one is used.
@@ 2.3 Implementation details
We conduct experiments with 18 MLLMs, including 11 open-weight MLLMs from the Llava-Next (Liu et al., 2024), Qwen2VL (Wang et al., 2024), Ovis-1.6 (Lu et al., 2024), and InternVL2.5 (Chen et al., 2024) families, ranging from 2 to 38 billion parameters. Additionally, we include four commercial models, GPT4 (OpenAI, 2023), GPT4o (OpenAI, 2023), Gemini-1.5-flash (Gemini-Team, 2024), and Gemini-1.5-pro (Gemini-Team, 2024), as well as three open-weight MLLMs specifically trained for visualization understanding: Llava-Chart-Instruct (Zeng et al., 2025), TinyChart (Zhang et al., 2024), and ChartGemma (Masry et al., 2025) ranging from 3 to 13 billion parameters. Qwen2.5-7B (Qwen-Team, 2024) serves as the text-only LLM. We use the transformers library (Wolf et al., 2020) to access all open models and the Azure API and Google Cloud API for GPT and Gemini models, respectively. We set the random seed at 42 and the temperature to 0. Following the standard ChartQA setup, all (M)LLMs are prompted in zero-shot. We report TinyChart results with the Direct approach (Zhang et al., 2024).
@@ 3.1 Experiment 1 - Assessing MLLM vulnerabilities (Accuracy)
Experiment 1 assesses the ability of MLLMs to answer questions about the underlying data given a visualization. The upper part of Figure 4 presents the QA accuracy of 18 MLLMs across the three datasets, sorted by increasing accuracy on ChartQA. The bottom part displays the accuracy for the three subsets constituting the misleading visualizations dataset. The results of Experiment 1 reveal three key findings.
@@
Performance is worse on misleading visualizations. All MLLMs perform substantially worse on misleading visualization compared to the non-misleading visualizations dataset, with accuracy dropping by up to 34.8 percentage points (pp), with an average of 23.4 pp. The decline is even more pronounced compared to ChartQA, reaching up to 65.5 pp, with an average of 54.0 pp. Furthermore, the mean MLLM accuracy on misleading visualizations (25.6%) is on par with the random baseline score (25.6%). These results demonstrate that MLLMs struggle to correctly interpret data distorded by misleaders.
@@
For comparison, average human accuracy reported on CALVI and CHARTOM is 80 and 89% on non-misleading visualizations, dropping to 39 and 49% on misleading ones, respectively (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024). On the same datasets, MLLMs achieve an average of 51 and 47% on non-misleading visualizations and 24 and 21% on misleading ones, falling short of human performance in all cases.
@@
The ranking of MLLMs by accuracy varies a lot across the three datasets. Notably, the two best-performing models on the CALVI subset of misleading visualizations rank among the worst on ChartQA, placing second and third from the bottom. This discrepancy suggests that achieving high accuracy on misleading visualizations is unlikely to emerge naturally from improvements on ChartQA, underscoring the need for dedicated mitigation methods.
@@
Performance is higher on real-world misleading visualizations than on other subsets. MLLMs perform better on the real-world subset of the misleading visualizations dataset, achieving an average accuracy of 33.5%, compared to 24% on CALVI and 21% on CHARTOM. In general, the best-performing MLLMs on misleading visualizations, GPT4o, Gemini-1.5-pro, and InternVL2.5-38B, primarily distinguish themselves through their higher performance on this subset. We hypothesize that this advantage stems from their parametric knowledge, which extends beyond 2022, the endpoint of the real-world subset. This allows these models to answer some questions based on stored knowledge about past events and statistics, without relying on the content of the visualization. In such cases, parametric knowledge effectively serves as a form of protection against real-world misleaders. However, this finding should be interpreted with caution. In practical scenarios, the underlying data will often be recent and therefore unlikely to be covered by the model’s parametric knowledge.
Figure 5: QA Accuracy (%) per type of misleader.
@@
Performance varies by type of misleader. Not all misleading visualizations are equally deceptive. This was already observed with human readers (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024), and is here further confirmed with MLLMs. Detailed accuracy scores per misleader and MLLM are reported in Figure 5. The three misleaders on which MLLMs perform best are concealed uncertainty, misleading annotations, and misrepresentation, with average accuracies of 64.8, 50.0, and 43.7, respectively. The four on which they perform worse are overplotting and inappropriate use of pie charts, both at 0.0%, and area encoding and missing data at 5.6%. For the two most represented misleaders, truncated axis and inverted axis, the MLLMs achieve average accuracies of 23.5 and 27.4 %, respectively. There are notable differences in which misleaders pose greater challenges to MLLMs compared to humans, and vice versa. For instance, area encoding is highly misleading for MLLMs, but not for human readers, who achieve an average accuracy of 91.5% on this type of misleader (Rho et al., 2023). Conversely, both humans and MLLMs struggle with misleaders such as overplotting, missing data, and 3D effects (Ge et al., 2023). The use of a dual axis is a misleader that poses greater difficulty for humans, who score an average accuracy of 16.1% on CHARTOM (Rho et al., 2023), compared to MLLMs, which achieve an average of 35.4%. These results highlight a notable difference in how humans and MLLMs process visualizations and are deceived by misleaders. This suggests that the methods needed to make MLLMs more robust to misleading visualizations will, at least in part, differ from those designed to protect human readers (Fan et al., 2022).

### v5

@@ Abstract
Visualizations play a pivotal role in daily communication in an increasingly data-driven world. Research on multimodal large language models (MLLMs) for automated chart understanding has accelerated massively, with steady improvements on standard benchmarks. However, for MLLMs to be reliable, they must be robust to misleading visualizations, i.e., charts that distort the underlying data, leading readers to draw inaccurate conclusions. Here, we uncover an important vulnerability: MLLM question-answering (QA) accuracy on misleading visualizations drops on average to the level of the random baseline. To address this, we provide the first comparison of six inference-time methods to improve QA performance on misleading visualizations, without compromising accuracy on non-misleading ones. We find that two methods, table-based QA and redrawing the visualization, are effective, with improvements of up to 19.6 percentage points. We make our code and data available.111github.com/UKPLab/arxiv2025-misleading-visualizations
Protecting multimodal large language models against
misleading visualizations
Jonathan Tonglet1,2,3, Tinne Tuytelaars2, Marie-Francine3 Moens Iryna Gurevych1 1 Ubiquitous Knowledge Processing Lab (UKP Lab), Department of Computer Science, TU Darmstadt and National Research Center for Applied Cybersecurity ATHENE 2 Department of Electrical Engineering, KU Leuven 3 Department of Computer Science, KU Leuven www.ukp.tu-darmstadt.de
@@ 1 Introduction
In an increasingly data-driven world, visualizations are widely used by scientists, journalists, governments, and companies to efficiently communicate data insights to a broad audience (Huang et al., 2025). They play a crucial role in crisis settings, such as during the COVID-19 pandemic, where charts were shared daily to inform the public (Zhang et al., 2021). In many cases, visualizations support a message more convincingly than if the underlying data table was shown directly to readers (Pandey et al., 2014). However, visualizations can also be misleading (Correll and Heer, 2017). This occurs when design flaws, or misleaders, distort the underlying data in a way that prevents a correct interpretation by the reader (Tufte and Graves-Morris, 1983; Pandey et al., 2015; Correll and Heer, 2017; Lauer and O’Brien, 2020; McNutt et al., 2020; Yang et al., 2021; Lo et al., 2022; Lisnic et al., 2023; Lan and Liu, 2025). These misleaders appear across a wide range of visualizations, including bar and line charts, or choropleth maps. Recent taxonomies have documented over 70 types of misleaders observed in real-world examples (Lo et al., 2022; Lan and Liu, 2025). Major categories include axes manipulations, such as truncated and inverted axes, or the use of visual manipulations like 3D effects. Figure 1 shows three real-world examples of misleading visualizations (Lo et al., 2022) paired with multiple choice questions (MCQs). Given a question about the underlying data table, their misleaders lead readers to infer a misleading answer instead of the correct one. Figure 2 illustrates how two visualizations of the same data, one non-misleading and the other misleading, can lead to different interpretations by readers.
Figure 1: Three examples of real-world misleading visualizations (Lo et al., 2022) with MCQs. The correct answer is colored in green, while the wrong answer supported by the misleader is colored in purple.
Figure 2: Non-misleading and misleading visualizations of the same data (Lauer and O’Brien, 2020) with a Likert-scale question where 1 means “a little” and 6 means “a lot”. A consistent interpretation requires identical responses. However, the deceived reader chooses a higher value for the truncated bar chart.
Prior work has revealed the potential of misleading visualizations to deceive human readers. Some studies (Pandey et al., 2015; Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024) have shown that such visualizations reduce human accuracy when answering MCQs about the underlying data. Other studies (Pandey et al., 2015; O’Brien and Lauer, 2018; Lauer and O’Brien, 2020) have demonstrated that human readers tend to provide different Likert-scale answers to a question depending on whether they view a misleading or non-misleading visualization of the same data table.
@@
Multimodal large language models (MLLMs) have sparked a massive interest in automated chart understanding research (Huang et al., 2025), with steady progress on reference benchmarks such as ChartQA (Masry et al., 2022; Hoque et al., 2022). However, existing research has largely overlooked the important threat posed by misleading visualizations. If MLLMs, like humans, are easily deceived by misleading visualizations, malicious actors could exploit this vulnerability. MLLMs confronted to misleading visualizations could present incorrect interpretations of the underlying data to users, contributing to the spread of disinformation, and reinforcing human belief in it. These risks underscore the urgency of thoroughly assessing the vulnerability of MLLMs to misleading visualizations and developing effective mitigation strategies.
In this work, we conduct the first comprehensive study to assess and mitigate the vulnerability of 19 MLLMs of varying sizes to misleading visualizations. In Experiment 1, we compare the question-answering (QA) accuracy of MLLMs on misleading and non-misleading visualizations. In Experiment 2, we evaluate the consistency of MLLMs, i.e., whether they provide the same response to a Likert-scale question when shown either a misleading or a non-misleading visualization of the same underlying data. The results of both experiments demonstrate that MLLMs, like humans, are indeed vulnerable to misleading visualizations. In Experiment 3, we compare six inference-time correction methods to reduce the vulnerability of MLLMs to the misleading visualizations of Experiment 1.
In summary, our contributions are twofold: (1) We present the first extensive analysis of the vulnerabilities of 19 MLLMs to 17 types of misleading visualizations, evaluating both QA accuracy and Likert-scale consistency. (2) We propose the first analysis of six correction methods to mitigate the negative impact of misleaders on QA tasks.
@@ 2.1 Datasets
Experiments 1 and 3 We compare the question-answering (QA) accuracy of MLLMs across three datasets: (a) a misleading visualization dataset containing $n$ = 143 instances, and featuring 17 types of misleaders, defined in Appendix A (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024; Lo et al., 2022); (b) a non-misleading visualization dataset ($n$ = 124) (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024; Lee et al., 2017), also combining real-world and synthetic cases; and (c) ChartQA (Masry et al., 2022), the reference real-world non-misleading benchmark ($n$ = 2500). Datasets (a) and (b) combine three existing resources and one introduced in this work. First, CALVI (Ge et al., 2023) includes 45 misleading and 15 non-misleading visualizations based on synthetic data, each paired with an MCQ. CALVI was initially designed and is the reference resource for evaluating humans. Second, CHARTOM (Rho et al., 2023; Bharti et al., 2024) contains 56 samples, including 28 MCQs, 20 free-text questions, and 8 rank questions. Each question is linked to two visualizations, one misleading and one non-misleading. Like CALVI, the underlying data is synthetic, and the test was originally designed to evaluate humans. Third, VLAT (Lee et al., 2017), the reference dataset to evaluate humans on non-misleading cases, provides 12 visualizations, each paired with three to seven MCQs, for a total of 53 instances. The visualizations are based on real-world data. Fourth, we introduce a dataset of 42 real-world misleading visualizations, each annotated with an MCQ with three to four choices. They come from a collection annotated with the misleaders affecting them (Lo et al., 2022), which inspired the synthetic examples in CALVI. We manually create MCQs, using as a reference those from CALVI and CHARTOM. By incorporating visualizations and questions about real-world data, we introduce direct conflicts with MLLMs’ parametric knowledge, allowing us to assess their vulnerability in real-world scenarios.
Experiment 2 We use four pairs of visualizations, one misleading and the other non-misleading, designed by Lauer and O’Brien (2020).
@@ 2.2 Correction methods
Providing (2) the axes data, (3) the data table, (4) or both: we prompt the MLLM in zero-shot to extract the axes labels or underlying data table. The axes and tables are formatted as text strings. We do not impose formatting constraints on the axes and tables. The axes, the table, or both are provided as additional prompt input.
(5) Table-based QA: after extracting the table with the MLLM, we provide it alone to a text-only LLM, reframing the task as table-based QA (Liu et al., 2023a, b; Kim et al., 2025).
(6) Redrawn visualization: after extracting the table with the MLLM, we provide it to a text-only LLM, which generates Python code to create a visualization using Matplotlib (Hunter, 2007). If the code compiles successfully, the generated visualization replaces the original one in the QA prompt; otherwise, the original one is used.
@@ 2.3 Implementation details
We conduct experiments with 19 MLLMs on a machine with two A100 GPUs, including 11 open-weight MLLMs from the Llava-Next (Liu et al., 2024), Qwen2VL (Wang et al., 2024), Ovis1.6 (Lu et al., 2024), and InternVL2.5 (Chen et al., 2024) families. Additionally, we include five commercial models, GPT4 and GPT4o (OpenAI, 2023), Gemini-1.5-flash and -pro (Gemini-Team, 2024), and Claude-3.5-sonnet (Anthropic, 2024), as well as three open-weight MLLMs specialized in chart understanding: Llava-Chart-Instruct (Zeng et al., 2025), TinyChart (Zhang et al., 2024), and ChartGemma (Masry et al., 2025). Qwen2.5-7B (Qwen-Team, 2024) serves as the LLM for table-based QA and visualization redrawing. We use the transformers library (Wolf et al., 2020) to access all open-weight models. All prompts are provided in Appendix B. We set the temperature to 0. Following the standard ChartQA evaluation setup, all (M)LLMs are prompted in zero-shot. We use the Direct approach for TinyChart (Zhang et al., 2024).
@@ 3.1 Experiment 1 - Assessing MLLM vulnerabilities (Accuracy)
Experiment 1 assesses the ability of MLLMs to answer questions about the underlying data given a visualization. The upper part of Figure 4 presents the QA accuracy of 19 MLLMs across the three datasets, sorted by increasing accuracy on ChartQA. The bottom part displays the accuracy for the three subsets constituting the misleading visualizations dataset. The results of Experiment 1 reveal three key findings.
@@
Performance is worse on misleading visualizations. All MLLMs perform substantially worse on misleading visualization compared to the non-misleading visualizations dataset, with accuracy dropping by up to 34.8 percentage points (pp), with an average of 23.8 pp. The decline is even more pronounced compared to ChartQA, reaching up to 65.5 pp, with an average of 53.8 pp. Furthermore, the mean MLLM accuracy on misleading visualizations (26.4%) is comparable to the random baseline score (25.6%). These results demonstrate that MLLMs struggle to correctly interpret data distorted by misleaders.
@@
For comparison, average human accuracy reported on CALVI and CHARTOM is 80 and 89% on non-misleading visualizations, dropping to 39 and 49% on misleading ones, respectively (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024). On the same datasets, MLLMs achieve an average of 52 and 48% on non-misleading visualizations and 24 and 22% on misleading ones, falling short of human performance in all cases.
@@
The ranking of MLLMs by accuracy varies a lot across the three datasets. The two best-performing models on the CALVI subset of misleading visualizations rank among the worst on ChartQA, placing second and third from the bottom. This suggests that achieving high accuracy on misleading visualizations is unlikely to emerge naturally from improvements on ChartQA, underscoring the need for dedicated mitigation methods.
@@
Performance is higher on real-world misleading visualizations than on other subsets. MLLMs perform better on the real-world subset, achieving an average accuracy of 34.7%, compared to 24% on CALVI and 22% on CHARTOM. In general, the best-performing MLLMs on misleading visualizations, Claude-3.5-sonnet, GPT4o, and Gemini-1.5-pro, primarily distinguish themselves through their higher performance on this subset. We hypothesize that this advantage stems from their parametric knowledge, which extends beyond 2022, the endpoint of the real-world subset. This allows these models to answer some questions based on stored knowledge about past events, without relying on the content of the visualization. Appendix C discusses further the impact of parametric knowledge.
@@
Performance varies by type of misleader. Not all misleading visualizations are equally deceptive. This was already observed with human readers (Ge et al., 2023; Rho et al., 2023; Bharti et al., 2024), and is here further confirmed with MLLMs. Detailed accuracy scores per misleader and MLLM are reported in Appendix D. The three misleaders on which MLLMs perform best are concealed uncertainty, misleading annotations, and misrepresentation, with average accuracies of 63.2, 50.5, and 44.4, respectively. The four on which they perform worse are overplotting and inappropriate use of pie charts, both at 0.0%, and missing data at 5.6%. For the two most represented misleaders, truncated axis and inverted axis, the MLLMs achieve average accuracies of 25.1 and 28.1 %, respectively. There are notable differences in which misleaders pose greater challenges to MLLMs compared to humans, and vice versa. For instance, area encoding is highly misleading for MLLMs, but not for human readers, who achieve an average accuracy of 91.5% on this type of misleader (Rho et al., 2023). Conversely, both humans and MLLMs struggle with misleaders such as overplotting, missing data, and 3D effects (Ge et al., 2023). The use of a dual axis is a misleader that poses greater difficulty for humans, who score an average accuracy of 16.1% on CHARTOM (Rho et al., 2023), compared to MLLMs, which achieve an average of 35.5%. These results highlight a notable difference in how humans and MLLMs are deceived by misleaders. This suggests that the methods needed to make MLLMs more robust to misleading visualizations will, at least in part, differ from those designed to protect human readers (Fan et al., 2022).